March BG Com Night 9, UT 2009 March 15

Observers: Dave Thompson (@LBT)
Telescope Operator: Aaron Ceranski
SW Support: Tony Edgin
Telescope Support: TShih, JRousseau

Summary:

  • 1 inch of snow up to sunset, cannot open dome
  • UPS overload problem, see IT#2067
  • SW Tests 1 and 3 passed.
  • Verified rotator center coordinates work as expected for RADEC and DETXY.

Details:

Sunset: clearing, but it had been snowing fairly heavily for the prior hour such that the enclosure has sufficient snow on top that we cannot open tonight.

~UT01:30 All computer room machines and user workstations suddenly shut down. It happened two more times as we tried to get things back up and running. See IT#2067 (critical) for details. It appears that the UPS serving the control room computers is overloaded and it tripped the circuit breaker #3 disconnecting the UPS from commercial power. The batteries drained until the UPS shut off power to connected computers. Since the computers (still) do not know when they are on UPS, they did not shut down cleanly. The load on the UPS dropped noticeably when we shut off and unplugged the four new TCS server machines in the computer room, as well as the LBC DVD burner. I recommend they stay off until the UPS overloading is resolved. We left the UPS plugged into commercial power but not connected to the computer to give the batteries a chance to charge a bit.

~UT06:30 we are back up with the remaining computers on the UPS circuit. We verified that we can send presets through iraf and everything seems to be working OK.

UT7:00 The LBC Red pump and cooldown appears to be good. T=147.6K, P=4.13E-06.

Note: The nature of the return message from the IRC commands has changed. They used to report back just a 0 or a 1 for success or failure, respectively. Now, the offset command returns "Resmsg[0]". Why was this changed? This is why the iraf scripts are reporting failure (because 0 = Resmsg[0]) even though the telescope is executing the requested command successfully.

UT10:19 TCS SW Tests: Test 1 - Evaluate the Offset Pointing functionality
pointto BS9153 (14:07:33.900 12:23:51.00) PA=0 Tel="TRACK"
verified that base RADEC are correct, offsets are zero, base DETXY are zero.

offset 14.65 15.00 type="ABS" coord="RADEC" (this is a 1.0 sec offset in RA)
Confirmed bases do not change, RADEC offsets do

zerobase RADEC
Confirmed, RADEC offsets are set to zero and the base RADEC is now at the offset position.

pointto BS9153 (14:07:33.900 12:23:51.00) PA=0 Tel="TRACK"
offset 16.67 16.67 type="ABS" coord="DETXY"
Confirmed only offsets in DETXY changed

zerobase DETXY
Confirmed offsets in DETXY are zeroed out, base in DETXY reflects the new position.

pointto BS9153 (14:07:33.900 12:23:51.00) PA=0 Tel="TRACK"
offset 14.65 15.00 type="ABS" coord="RADEC" (this is a 1.0 sec offset in RA)
offset 16.67 16.67 type="ABS" coord="DETXY"
Confirmed that both of the offsets are in effect, bases remained unchanged.

zerobase RADEC
Confirmed existing DETXY offsets were unchanged

offset 14.65 15.00 type="REL" coord="RADEC"
zerobase DETXY
Confirmed existing RADEC offsets were unchanged.

offset 14.65 15.00 type="REL" coord="RADEC"
Confirmed REL RADEC offsets behave as expected.

offset 16.67 16.67 type="REL" coord="DETXY"
offset 16.67 16.67 type="REL" coord="DETXY"
Confirmed REL DETXY offsets behave as expected.

Test 1 passed qualitatively. Need to remeasure on-sky for accuracy once the AGW coordinate system is calibrated.

UT11:05 TCS SW Tests: Test 2 - Evaluate the update Pointing Reference functionality
Test 3 passed. Only the requested coordinates are updated, the other remains unchanged.

Also checked that the new rotator center coordinates behave as expected. They do in RADEC. In DETXY space it is a little harder to interpret because of the Xflip. Currently, PCS's DETXY at PA=0 space has N-up and E-left even though LUCIFER has E-right. So, an offset of +29.3 arcsec in the +x DETXY direction moved the PO to the west and thus the rotator center to the east. Confirmed - increase of 2 seconds of time in RA. If the PO is moved to the north, then the rotator center moves south. Also confirmed.

Test 11 - confirmed that there is still a problem. DETXY offsets move the achieved coordinates by twice that amount. RADEC offsets work as expected.

UT11:26 Parking the telescope at horizon so Tim can pump LBC Blue in the morning.

-- DavidThompson - 15 Mar 2009
Topic revision: r3 - 27 May 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback